You have a square and a rectangle and are told their combined area is 67. You are told the rectangle has a width of 3 cm more than the length of a side of the square and the rectangle has a width 2 cm more than the length of a side of the square. Find the dimensions of the rectangle.

Answer by edjones(8007)

Let the side of the square equal s.
s^2+(s+3)(s+2)=67
s^2+s^2+5s+6=67
2s^2+5s-61=0
s=(3sqrt(57)-5)/4) Quadratic formula
(s+3)(s+2)
=(((3sqrt(57)-5)/4)+3)cm. and (((3sqrt(57)-5)/4)+2) cm. The dimensions of the rectangle.
.
Ed